An internet address is 32 bits long. The first 8 bits are dedicated to the netid and the rest to the hostid. The netid must start with a "0" and cannot include the address"01111111" THe hostid made up of all 0's or all 1's are not allowed using this scheme how many different internet addresses are available for computers on the internet.
Answer #1

Let's see the number of network possible by considering number of possible netId's.

In 8-bits , first bit is 0 and remaining 7 bits can be either 0 or 1.so,for remaining 7 bits can have 2 bits.so,total number of possible netId's = 27

But,it is given that 01111111 can't be netId so,

Total number of netId's possible = 27-1 = 127

Now,for hostID,there are 24 bits in host portion.

For each bits there is two possibility i.e 0 or 1 .so,total number of possible hostIds = 224

But two hostIds(all 0's and all 1's) are not to be included so,total number of hostIDs possible = 224-2

total number of IP addresses available for computers on internet = total number of netIds * total number of hostIds

= 127 * (224-2)

= 127 * 16777214

= 2130706178